Circular No. 8329

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004bj AND 2004bk Further to IAUC 8321, H. Pugh and W. Li report the LOSS discovery

of two apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004bj Apr. 22.36 13^{h}16^{m}18.^{s}03 + 7^{o}02′52′′.7 18.3 15′′.0 E, 6′′.3 N
2004bk Apr. 22.36 13 37 28.72 + 4 06 11.3 16.0 9′′.9 W, 4′′.8 S

Additional KAIT magnitudes: SN 2004bj in MCG +01-34-13, Mar. 21.35 UT, [19.0; Apr. 2.38, [18.0; 23.35, 18.3. SN 2004bk in NGC 5246, Apr. 2.38, [19.0; 23.35, 15.9.

COMET C/2002 T7 (LINEAR) E. S. Howell, Arecibo Observatory, reports that observations of the

18-cm OH lines in this comet, made with A. J. Lovell and F. P. Schloerb at Arecibo Observatory, result in the following production rates: Apr. 10.63 UT, log Q(OH) = 29.4; 15.60, 29.3; 16.59, 29.4; 17.59, 29.5 (uncertainties of order 10^{28} mol/s). Line shapes are consistent with parent outflow velocities ranging from 1.7 to 2.0 km/s. Spectra obtained on Apr. 14–15 reveal an excess of 10–20 percent around cometocentric velocities of +3 km/s.

1999 DJ_{4}

L. A. M. Benner, Jet Propulsion Laboratory; M. C. Nolan, National Astronomy and Ionospheric Center, and Arecibo Observatory; S. J. Ostro and J. D. Giorgini, Jet Propulsion Laboratory; J. L. Margot, University of California, Los Angeles; and C. Magri, University of Maine at Farmington, report: “Arecibo delay-Doppler images (2380 MHz, 12.6 cm), obtained on Apr. 14–16, confirm that 1999 DJ_{4} is a binary system (cf. IAUC 8316). Preliminary estimates of average diameters, based on range extents at 15m resolution, are 420 and 200 m. The maximum range separation observed is more than 700 m.”

COMET C/2004 F4 (BRADFIELD) Visual total-magnitude estimates in bright twilight: Apr. 22.48 UT,

4.5 (A. Hale, Cloudcroft, NM, 0.2-m reflector); 23.21, 3.9 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector; 0^{o}.1 tail in p.a. 20^{o}); 23.48, 4.5: (Hale, 10×50 binoculars; 15′ tail in p.a. 295^{o}).

2004 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green